OpenStudy (anonymous):

what is the equation solutions x3-3x+27=8x-3

hero (hero):

subtract 8x - 3 from both sides to get: x^3 -11x 30 = 0

jhonyy9 (jhonyy9):

x^3 -11x +30 =0 here is right

jhonyy9 (jhonyy9):

- so but i think that you want writing the first term x^2 and if this is true your equation can be writing in the form (x-5)(x-6)=0 solutions are 5 and 6 x_1=5 x_2=6